OpenStudy (aizhalee):

Fan and Medal will be given ! In order to build a scale model of the trail, the drawing is enlarged on the coordinate plane. If two corners of the trail are at point A(−2, 7) and point D(−10, −1), what is another point that could represent a corner of the trail?

OpenStudy (aizhalee):

theres a big test coming next week and I need help working on this question , can someone please help me with questions like these ?

OpenStudy (sdfgsdfgs):

u have picked the right ans A :)

OpenStudy (aizhalee):

i guessed tho , I want to know how to get it @sdfgsdfgs

OpenStudy (aizhalee):

Can you help me ?

OpenStudy (sdfgsdfgs):

OK all 4 choices have 7 as the y-coordinate. so no need to worry about that right?

OpenStudy (aizhalee):

Yes .

OpenStudy (sdfgsdfgs):

look at A and D, what is the difference between their y-coordinates? then look at G and H, what is the difference there?

OpenStudy (aizhalee):

A y coordinate is 7 and D ycoordinate is -1 . G y coordinate is -2 and H y coordinate is -4

OpenStudy (sdfgsdfgs):

So the difference between AD is 7-(-1)=8 GH is -2-(-4)=2 what is the ratio of the two differences?

OpenStudy (aizhalee):

8:2 ? simplified 4:1 ?

OpenStudy (sdfgsdfgs):

Yup. so u now know the ratio between the new and old shape: it is 4:1

OpenStudy (aizhalee):

Thank you so much so how do I use that to find the other end point of the parralelogram ?

OpenStudy (sdfgsdfgs):

What is the difference in X-coordinates between F and G?

OpenStudy (aizhalee):

4 and 7

OpenStudy (sdfgsdfgs):

n the difference is....

OpenStudy (sdfgsdfgs):

difference is to subtract one from the other...

OpenStudy (aizhalee):

7-4=3 **

OpenStudy (sdfgsdfgs):

Good :) with the ratio, what should be the difference in the bigger shape? Hint: it is 3*something...

OpenStudy (aizhalee):

with the ratio being 4:1 and the small shape is 3 spaces apart , the bigger shape has to be 12 ?

OpenStudy (aizhalee):

4 times 3 ?

OpenStudy (sdfgsdfgs):

Great job! Now because that is the difference of the larger shape, all u have to do is adding that back to the X-coordinate of A so -2 + (4 times 3) =?

OpenStudy (aizhalee):

im writing these in my notes

OpenStudy (sdfgsdfgs):

it is -2 + 12...

OpenStudy (aizhalee):

sorry with the parenthesis around the 4 +3 i thought i had to multiply

OpenStudy (aizhalee):

-2 + 12 = 10

OpenStudy (sdfgsdfgs):

BINGO n thats the ans u guessed correctly lol

OpenStudy (aizhalee):

Thank you sooooooooooooooooooo much !♥ I could hug you right now . I appreciate it :D !

OpenStudy (sdfgsdfgs):

welcome n hug u back @Aizhalee